OpenStudy (seniaa.tapia):

simplifying rational exponents (216r^9)^1/3

OpenStudy (jdoe0001):

@seniaa.tapia so.. if you factor 216, what factors can it give you?

OpenStudy (anonymous):

216=6*6*6=6^3